An economy produces three goods: orange, apple and orange juice. It is assumed that half of the orange are bought and consumed as final good; the other half is used to produce orange juice.                                              

In 2019, 200 KGs of oranges are produced at Rs.40/kg, 500 KGs of apples are produced at 120/kg, and 200 Glasses of orange juice are produced at Rs.150/glass.

In 2020, 220 KGs of oranges are produced at Rs.50/kg, 450 KGs of apples are produced at 140/kg, and 230 glasses of orange juice are produced at Rs.135/glass.

Assuming 2019 is the base year, calculate the GDP Deflator in 2020.



1
Expert's answer
2021-05-05T13:36:11-0400

Solution:

GDP Deflator = "\frac{Nominal

\;GDP}{Real \;GDP} \times 100"


Therefore, we first, calculate the values for Nominal GDP and Real GDP

 

Nominal GDP = Sum of (this year prices "\times" this year quantities)


GDP only accounts for final goods, therefore we will only take half of the oranges produced and bought and consumed as a final good. The other half is an intermediate good used to produce orange juice and hence will not be added to the GDP to avoid duplication.

 

Nominal GDP 2020:

Oranges = 50 "\times" 110 = 5,500

Apples = 140 "\times" 450 = 63,000

Orange juice = 135 "\times" 230 = 31,050

Total = 5,500 + 63,000 + 31,050 = 99,550

Nominal GDP 2020 = 99,550

 

Real GDP = Sum of (base years prices "\times" this year quantities)

 

Real GDP 2020:

Oranges = 40 "\times" 110 = 4,400

Apples = 120 "\times" 450 = 54,000

Orange juice = 150 "\times" 230 = 34,500

Total = 4,400 + 54,000 + 34,500 = 92,900

Real GDP 2020 = 92,900

 

GDP Deflator = "\frac{Nominal

\;GDP}{Real \;GDP} \times 100"


GDP Deflator = "\frac{99,550}{92,900} \times 100 = 107.16\%"


GDP Deflator year 2020 = "107.16\%"